OpenStudy (anonymous):

Part 1: Using complete sentences, explain how to find the equation of the line, in standard form and slope–intercept form, passing through (3, 6) and (–2, –4). (4 points) Part 2: Compare the benefits of writing an equation in standard form to the benefits of writing an equation in slope–intercept form. (3 points)

OpenStudy (anonymous):

Line equation through two points A(x1,y1) and B(x2,y2) is given by y-y1=(y2-y1)/(x2-x1) (x-x1) in your case y-6=(-4-6)/(-2-3) (x-3) y-6=2(x-3) y=6+2x-6 y=2x

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

yes i am can you please explain with words because i dont understand

OpenStudy (anonymous):

It is all to do with linear equation. so, slope–intercept form. is better cause u see slope and intersept point and in standard form all is mixed up

OpenStudy (anonymous):

i understand that i dont understand the first part

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

In order to write an equation of a line, you need a point (intercept) and a slope. You already have a good point to use (3,6) in your slope-intercept form -> y - y = m ( x - x ) -> y - 6 = m ( x - 3 ) Now all you need is the slope. To find the slope you need to (y2-y1) / (x2-x1). So that's [ 6 - (-4) ] / [ 3 - (-2) ] = 10 / 5 = 2 m= 2 so you can add it into your equation. y - 6 = (2) ( x - 3 ) To put it into standard form (y=mx + b) you must solve for y. y - 6 = 2x - 6 y= 2x

OpenStudy (anonymous):

i just had a epiphany thankyou